# An Adaptive Dual-Domain Prediction Strategy based on Second-order Derivatives for Dynamic Multi-Objective Optimization

### v1

@@ IV-D2 Ablation Study
TABLE V: COMPARISON OF RESULTS OF ADPS-DMOEA VERSUS ITS TWO VARIANTS ON DF TEST PROBLEMS
@@
Table V illustrates the results from an ablation study designed to evaluate the efficacy of ADPS-MOEA/D and its two single-domain variants when faced with DF test problems. The symbols ”+”,”-”and ”=” denote our proposed ADPS-MOEA/D being statistically significantly better, worse, or equivalent to the ablation algorithms, respectively. This analysis clearly demonstrates the superior performance of the dual-domain ADPS-MOEA/D, which operates in both the decision and objective spaces, emphasizing the value of our adaptive dual-domain approach. The adaptive approach allows the algorithm to effectively combine insights from both decision and objective spaces, while dynamically weighting the contributions from each domain in response to dynamically changing environments. This achieves significantly more reliable and superior performance outcomes than either of the single-domain variants. This capability is particularly critical in dynamic environments where conditions and requirements can shift unpredictably.
@@
Fig. 6 illustrates the performance of the ADPS algorithm alongside its variants, ADPS-I and ADPS-II, across 30 environmental changes. The left graph displays the IGD for the DF1 benchmark, while the right graph shows the IGD metric for the DF13 benchmark. The plots reveal that ADPS experiences a notable increase in IGD at the outset, indicating a significant initial change, to which ADPS adapts more effectively than its variants. Throughout the environmental changes, ADPS shows lower IGD values, signifying its superior ability to maintain solution quality across dynamically changing scenarios. In contrast, ADPS-I and ADPS-II exhibit higher and more fluctuating IGD peaks (especially clear in Fig. 6 right-hand graph for problem DF13), suggesting reduced stability and adaptability in response to environmental shifts. The large periodic spikes observed in the ADPS-I variant (based in the decision domain), illustrate how certain types of environmental changes can pose greater challenges for variants utilising one domain versus another. This underscores the greater robustness provided by the full ADPS algorithm, which: i) exploits information from both domains; and ii) can adaptively weight its attention towards one domain or the other, as the environmental circumstances change over time.
@@
Figure 7: Averaged MIGD values obtained by ADPS-MOEA/D with different adaptive adjustment step sizes during dual-domain prediction on DF1, DF5, DF9, and DF13.
IV-D3 Sensitive Study of Parameter $\lambda$ in dual-domain prediction
In our proposed method, the size of the training set for each domain is equal, but the adaptive adjustment step size for different domains may affect the performance of the prediction approach. To study the sensitivity of dual-domain prediction to the adaptive adjustment step size, we set a hyper-parameter $\lambda$. This represents the step size of adaptive adjustment in the weighting of the search space within the entire initial population. Specifically, we tested five different values of $\lambda$ (ranging from 0.01 to 0.05), on four representative benchmark problems (DF1, DF5, DF9, and DF13), which are chosen because they exhibit significantly different change patterns in the decision and objective spaces. Furthermore, four different configurations ($n_{t}$ = 1, 2.5, 5, 10, $\tau_{t}$ = 10) were chosen to evaluate the performance of our method in terms of adaptive adjustment of weighting between the two domains. A parameter value exceeding 0.05 surpasses the population size during 30 environmental changes, rendering it insignificant. Figure 7 shows the averaged MIGD values obtained by our method on these test instances. The results indicate that our method performs relatively poorly when the value of $\lambda$ is very large (i.e., $\lambda$ = 0.04, 0.05), while the algorithm performs significantly better on these benchmarks when $\lambda$ is assigned a value close to 0.02. From these results, we can infer that prudent selection of the step size $\lambda$ for adaptive adjustment in domain weighting, is an important consideration for optimal prediction of dynamic changes when solving DMOPs.
IV-E Summary of Results
In this section, we have presented the results of two sets of experiments, where we evaluated the performance of four advanced and classic methods against the proposed ADPS-MOEA/D strategy, using aqn extensive set of benchmark test problems. Fourteen test instances were used to assess the performance of ADPS-MOEA/D. According to the results presented, the ADPS-MOEA/D method proposed in this paper exhibits superior predictive capabilities, and can find solutions with better diversity and distribution, compared to the compared state-of-the-art methods from the literature. Furthermore, ablation experiments clearly demonstrated the advantages of ADPS in its ability to adaptively combine operations in both the decision and objective domains, as compared to ablted methods operating in either domain alone.

### v2

@@ IV-D2 Ablation Study
TABLE IV: COMPARISON OF RESULTS OF ADPS-DMOEA VERSUS ITS TWO VARIANTS ON DF TEST PROBLEMS
@@
Table IV illustrates the results from an ablation study designed to evaluate the efficacy of ADPS-MOEA/D and its two single-domain variants when faced with DF test problems. The symbols ”+”,”-”and ”=” denote our proposed ADPS-MOEA/D being statistically significantly better, worse, or equivalent to the ablation algorithms, respectively. This analysis clearly demonstrates the superior performance of the dual-domain ADPS-MOEA/D, which operates in both the decision and objective spaces, emphasizing the value of our adaptive dual-domain approach. The adaptive approach allows the algorithm to effectively combine insights from both decision and objective spaces, while dynamically weighting the contributions from each domain in response to dynamically changing environments. This achieves significantly more reliable and superior performance outcomes than either of the single-domain variants. This capability is particularly critical in dynamic environments where conditions and requirements can shift unpredictably.